Pmg agenda feb 3 2000
Download
1 / 17

PMG Agenda Feb 3, 2000 - PowerPoint PPT Presentation


  • 131 Views
  • Uploaded on

PMG Agenda Feb 3, 2000. Overview Project Manager’s Report – Tuts Fallback options – Weerts Silicon Support Cylinder Modifications - Cooper Fallback Options & Installation - Kotcher. Completed Reportable Milestones. 3 early 3 on schedule (  |.2w|) 7 late. Near Term Reportable Milestones.

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about 'PMG Agenda Feb 3, 2000' - bernad


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
Pmg agenda feb 3 2000 l.jpg
PMG AgendaFeb 3, 2000

  • Overview

    • Project Manager’s Report – Tuts

    • Fallback options – Weerts

  • Silicon Support Cylinder Modifications - Cooper

  • Fallback Options & Installation - Kotcher


Completed reportable milestones l.jpg
Completed Reportable Milestones

  • 3 early

  • 3 on schedule (|.2w|)

  • 7 late



Status overview since last pmg l.jpg
Status Overview since last PMG

  • Silicon sensors

    • Axial (3 chip)

      • No problems

    • 90deg (6 chip)

      • Micron delivery an issue (sputterer problem at Micron, being fixed)

      • Micron delivery paces

      • Fallback – accept lower grade detectors

    • 2deg (9chip)

      • Production going well (last batch 19/25 passed)

      • Full delivery in April

    • F wedges

      • Proceeding well at 50-60% yield at Micron (last batch 5/6 passed)

      • Received 5/75 sensors from Eurisys

    • H wedges

      • 60 more on order


Silicon l.jpg
Silicon

  • HDI

    • Getting sufficient HDI’s to keep pace with production is a problem

      • Expediting additional HDI orders

      • 2nd vendor for bonding, Silitronics, looking good

    • H wedge

      • 2 HDI left, as expected

      • Production stops for 4 wks

    • 2deg (9chip)

      • Out of HDI’s, rounding up what is available 5-9

    • F wedge

      • Almost out of HDI’s


Silicon6 l.jpg
Silicon

  • Ladder & Wedge production

    • Axial (3 chip)

      • Picking up next week

    • 2deg (9chip)

      • Production lagging, out of HDI’s

    • 90deg (6chip)

      • Paced by Micron delivery

    • F wedges

      • Limited by HDI availability

    • H wedges

      • Limited by HDI availability



Fiber status l.jpg
Fiber Status

  • Fibers

    • Waveguide production going well

      • 175 bundles pulled & cut

    • Cylinders 8, 7 complete

    • Nesting cart is being measured; delays by about 4 weeks

      • 8a and 7a installed and glued

      • Final stiffness tests in progress

      • Start assembly next week, 2/7

    • Flex circuits still a concern, delaying complete start of VLPC cassette production

      • Compunetics 100pc order nearly complete – 34 good, 31 to go

      • 1st production cassette under test

      • Many components done while awaiting flex circuits (~20%)


Fiber tracker status l.jpg
Fiber Tracker Status

  • Ribbon Production

    • Continues


Fiber tracker status10 l.jpg
Fiber Tracker Status

  • Ribbon Mounting

    • Continues

      • C8 and C7 Done

      • C6 axial started


Status overview l.jpg
Status Overview

  • Preshowers

    • Some schedule re-arrangement, net speedup

  • Tracking Electronics

    • Some (2-4wks) slippage in layout and vendor production, worry here…….

  • Calorimeter Electronics

    • 98% of preamps (59kch) in hand

    • Power supplies complete

    • 80% of packaged SCA’s in hand, will need to order 5k more

    • Small layout problems cause 6 week delay to BLS schedule

    • Installed 2 power supplies & 16 card of preamps in detector with online controls for p.s.


Status overview12 l.jpg
Status Overview

  • ICD

    • Module testing underway

    • Preamps in hand, motherboard ordered

  • Central Muons

    • Commissioning continues

    • First ½ truss in D0 Hall

  • Forward Muons

    • Trig Pixels: Proceeding well on schedule.

    • See no problems meeting April completion date

    • MDTs: @Dubna completed. Being shipped in heated container.

    • Production on A MDTs layer has started

    • C MDT parts presently most critical items (panels, etc)



Trigger pixels l.jpg
Trigger pixels January 31, 2000

The forward muon trigger planes:

3 planes: A, B, C with 16 octants ( north & south) each

All pixel counters are at FNAL

Layers A & C are completed

Layer B started


Mdt chambers octant assembly l.jpg
MDT chambers & octant assembly January 31, 2000

The tracking planes:

MDT production at Dubna complete ( shipping)

A layer assembly done


Status overview16 l.jpg
Status Overview January 31, 2000

  • Trigger & DAQ

    • Proceeding well

    • Delays in some in 12MCM boards ( stereo CFT, CPS, FPS), because of engineering

  • Online

    • Manpower in good shape

    • No schedule concerns, not on critical path

    • Being tapped for manpower in other systems


Summary l.jpg
Summary January 31, 2000

  • Concerns

    • HDI deliveries

    • Silicon testing

    • Fiber flex circuits for VLPCs

    • Forward muon installation

    • M&S cost contingency continues to dwindle

  • On the horizon

    • Lehman review…….


ad